# Characterizing variability in complex network community structure with a recursive significance clustering scheme

### v1

@@ Abstract
Network science has presented community detection as a valuable tool for revealing the functional modules in complex systems as rooted in the wiring architectures of complex networks. The varying procedures of community detection can produce, however, divisions of a network into communities that are prone to degeneracy given small changes to the network’s configuration. This yields network partitions of similar merit that are structurally dissimilar, which is especially problematic when the network is constructed on uncertain data. To reconcile with the ambiguity in interpreting degenerate network partitions as representations of underlying system function, we introduce a recursive significance clustering scheme that identifies the subsets of nodes that have stable joint community assignments under network perturbation. These robust node groups are referred to here as cores, and represent well-supported features of the network. We show that cores characterize the variability inherent to non-overlapping community structure in weighted and directed networks and indicate robust community assignments that are cohesive under temporal evolution of the network.
@@ I Introduction
Characterizing complex systems is one of the great modern scientific challenges owing to the difficulty in accurately and completely deriving their collective behaviours from their atomic components [1]. Enter complex networks, whose anatomies encode the intricate interactions between the discrete components of a complex system [2, 3]. Despite varying greatly in form, composition, and scope, complex networks often share common organizing principles [4].
Of particular interest in network science is the structure of a network’s wiring architecture, since structure of the network represents function of the system [1]. One manner in which many complex networks arrange themselves is into communities: groups of nodes with high internal link density that are sparsely linked with others [5]. In non-overlapping community structure, each node is assigned to exactly one community. Communities may be understood as representations of functional modules embedded within the networked system [6]. Scaling the lens of analysis to the mesoscopic interactions among communities omits superfluous detail without obscuring significant relationships in the system, providing a simplified view of interaction patterns [7].
The problem of community detection, however, is ill-posed: there is no single definition of what constitutes a community that is universally accepted [6], and there is no free lunch [8], implying that there is no single community detection algorithm that is optimal for all applications. As community detection is an NP-hard problem [9], most algorithms utilize stochastic searches [10, 11, 12, 13] to optimize an objective function that idiosyncratically defines the quality of a network partition [14, 15]. This exposes results to degenerate solutions of dissimilar community structure that are of similar merit [16, 17, 18, 19, 20]. Small changes to the network’s components (due to, e.g., measurement noise, model uncertainty, or temporal evolution) can exacerbate this issue, causing radically different structures to be detected [19, 21, 22].
@@
If a network exhibiting degenerate community structure makes the interpretation of a single partition misleading [16, 31], and consensus solutions may blur relevant information captured by different partitions [21], how then can the results of community analyses be interpreted in a scientific context? We now describe and demonstrate a recursive significance clustering scheme to reveal subsets of nodes, referred to here as cores, that are jointly assigned to a community within an ensemble of network partitions with significance (e.g., fig. 1c). A characterization of cores provides insight into the variability of community structure detected in an empirical network, and especially one with inherent uncertainty, by indicating the community assignments that are well-supported by the data.
@@ II Recursive significance clustering
For the results of community detection to be of value, the extent to which detected communities are statistically significant as opposed to coincidence must be considered [27]. After all, it is possible that a network has very weak or no community structure, meaning that detected communities are fragile and merely an artifact of the detection method used [16, 21]. A distinct optimal solution should be robust to small changes to the network’s configuration owing to the gap in quality to competing optima. In turn, the matter of community structure significance is reformulated as a question of how robust that structure is to network perturbations [27].
To identify structure persistent across an ensemble of partitions, significance clustering procedures seek to distinguish between stable and unstable node assignments to communities [32, 21]. Stable or significant nodes of a community are flagged based on the proportion of ensemble partitions in which they are jointly assigned to a community. This approach discovers only the largest significant subset of nodes belonging to a community, but there may exist other mutually exclusive subsets of nodes in that community that are jointly assigned across the solution ensemble, but not with the largest or other subsets. This limitation betrays another: a single partition of the network is used as a reference with which to identify these significant subsets, necessarily dictating the quantity and delineation of significant subsets found. A single partition neglects relevant information available in a degenerate solution landscape [21], making this reliance on a reference partition as the mould for significant subsets undesirable.
We define the robust cores of a network to be subsets of nodes that are jointly assigned across the partition ensemble. For significance level $\alpha$, we require that all nodes of a core are jointly assigned in a $(1-\alpha)$-fraction of partitions to relax the requirement. The arrangement of cores into communities may vary among solutions, but that of nodes into cores is robust; in this light, cores are nearly-indivisible constituents of community structure. Then, delineating cores is akin to assessing the community structure variability, as they are the cohesive units that compose dissimilar partitions. Nodes that do not regularly cluster with any other nodes, and nodes belonging to cores too small relative to the size of the network to contribute meaningful insight into system function, are regarded as unstable or noisy assignments.
@@
    The empirical network has its components perturbed repeatedly to generate many replicate networks that together capture modelled variability of and/or uncertainty in the network’s configuration.
@@
    An optimization technique is used to find the largest subset of nodes that are assigned to the same community in a $(1-\alpha)$-fraction of partitions; then the next largest core is searched for, and so on recursively until all cores of meaningful size are extracted.
This recursive approach has the advantage of operating without use of a reference partition; instead, the solution ensemble is used intrinsically to discover the well-supported features of the network. Section VI provides a detailed account of how we implement each step.
@@ III Viewing community structure variability
We have established that cores are the nearly-indivisible groups of nodes that arrange themselves into communities across an ensemble of network partitions. Having knowledge of which cores unite with each other to form communities, and which stand alone, gives a glimpse into the structural patterns present in the partition ensemble. Bear in mind that this excludes the unstable (noisy) assignments that have been filtered or other nodes which on occasion are not jointly assigned with the rest of their core.
For this, we define the coalescence frequency of any combination of cores to be the proportion of replicate networks in which all of the nodes of those cores are assigned to the same community and with no other cores. Further, the stability of a core is the proportion of replicate networks in which the nodes of the core are jointly assigned, possibly with other cores. It is guaranteed by definition to be within the interval $[1-\alpha,1]$ for significance level $\alpha$. The sum of every coalescence frequency involving a core will equal that core’s stability.
Notably, the coalescence frequency of just one core gives the proportion of partitions where that core forms a stand-alone community. Cores with this self-coalescence frequency close to their stability are well-supported functional modules of the system. When this is the case for an entire set of cores, it signals very strong community structure in the network.
@@ IV Application
To illustrate this approach, we analyze cores of a real network. We begin by introducing the network and recursive significance clustering setup in section IV.1. We then look at the network cores and how they compose a community structure ensemble in section IV.2, and their temporal evolution in section IV.3. Finally, we briefly examine the nodes with unstable community assignments in section IV.4.
@@ IV.1 Alaska–Alaska flight network
Flight networks benefit from community detection to uncover the structure of flight routes [33, 34]. The network of flights in Alaska (AK), United States serves as an interesting case study thanks to the key role flight plays in the region’s transportation system [35].
We consider three temporal snapshots of the AK–AK flight network: calendar years 2017, 2020, and 2023. In the 2017 (resp. 2020; 2023) network, there are 3,578 (3,285; 3,616) directed links weighted by the number of aggregate flights in that year from one of the 334 (328; 330) airports to another, for 401,153 (275,890; 364,285) total annual flights. (See section VI.1 for information on data compilation). These years were selected to capture flight traffic before, during, and after the largest impacts of the COVID-19 pandemic, an unparalleled disruption to aviation activity [36] that admits a prime opportunity to analyze the temporal behaviour of robust network cores.
We assume flights are independent events and resample each link’s weight from a Poisson distribution with rate parameter equal to that link’s weight in the empirical network to form 1000 replicate networks [32]. We then find the robust cores of the network’s community structure, detected using Infomap [12], at a significance level of 0.05, accepting only cores of at least five nodes. This is completed for each temporal snapshot. (See section VI for a description of this procedure in complete detail.)
@@ IV.2 Reckoning with dissimilar structure
Figure 2: Cores of the 2017 temporal snapshot of the AK–AK flight network. (a) Spatially-embedded cores. Nodes with unstable community assignments are shaded grey. (b) Core coalescence frequency and stability.
Spatially-embedded cores of the 2017 AK–AK flight network snapshot with their coalescence frequencies and stability are shown in fig. 2 (find that of the 2020 and 2023 snapshots in appendix A). For visualization, we appropriate slightly the UpSet plot [37] to show the coalescence frequency for any combination of cores in tandem with the stability of each core. To manage possible combinatorial explosion, only core combinations with a coalescence frequency of at least 0.05 are displayed.
It is now simple to characterize the dynamics of community assignments across the resampled networks. For instance, cores 3, 4, 5, 6, 7, and 9 frequently (or always) form communities by themselves with no other cores. Oppositely, adjacent cores 2 and 8 almost always merge together as one community (with others) — on the cusp of their union being defined as a single core. This behaviour contrasts adjacent cores 3 and 5 in the east which never merge.
Speaking to the system, core 1 includes the Ted Stevens Anchorage International Airport (IATA: ANC) which brokers a tremendous amount of AK air traffic, making it a highly influential node in the network as measured by any centrality index. This is responsible for the vast geographic distribution of core 1 compared to other cores: the connectivity of ANC recruits airports to core 1 that otherwise have no strong local connection structure. The coalescence frequencies involving core 1 show that it sometimes joins the merging of cores 2 and 8, occasionally bringing core 6 with it, but also stands on its own or with core 4 or 6. We imagine this as it irregularly absorbing a few of the smaller cores in close proximity to it. This indicates some presence of rather coarse community structure engulfing the west coast (except core 7) and much of the inland in the ensemble.
@@ IV.3 Temporal cohesion of cores
Now we examine the cores of successive network snapshots to assess structural changes of the network in time. An alluvial diagram [32] shows this in fig. 3. Streamlines show the core assignments of nodes changing between snapshots, coloured by their initial core assignment to track them in time. Since the node set of the network is not static in time, the flow of a node is not shown if it was not present in a contiguous temporal snapshot.
Figure 3: Core assignment changes of nodes between the 2017, 2020, and 2023 temporal snapshots of the AK–AK flight network. Nodes with unstable community assignments are placed in the grey-shaded block.
It is immediately apparent that each core largely remains cohesive in time, i.e., subsets of nodes continue to be jointly assigned through successive temporal snapshots. The temporal evolution of cores in the network, even over a period containing a major disruption to flight volume, is largely understood by the initial set of cores, their coalescence frequencies, and stabilities. We note a few details supporting this in particular:
  • Core 1, which has relatively low stability in 2017, splits, forming a new, small core in 2020 that later merges back in 2023. Some of its other nodes become unstable.
  • Cores 2 and 8, on the borderline of being a unitary core in 2017, merge together cleanly in 2020 and through to 2023.
  • Cores 5, 6, 7, and 9 persist through time with complete cohesion; likewise for cores 3 and 4, with few of their nodes becoming unstable.
The fusion of core 9 with core 1, while not anticipated by its stability or self-coalescence frequency, is not surprising either owing to the influence of ANC in the network. An anomalous change in flight volume between an individual airport and ANC can obfuscate or reveal small-scale substructure. An increase in flight volume from 2017 occurred between ANC and Beluga Airport (BVU) of core 9 to catalyze the merge with core 1 in 2020, and a decrease between ANC and Homer Airport (HOM) for the emergence of the smaller core split from core 1. As these cores are relatively small, the core size threshold could be adjusted post hoc to exclude them.
@@ IV.4 Speculating on the nature of unstable nodes
Figure 4: Wiring architectures and strength distributions of nodes with unstable community assignments for the 2017 AK–AK flight network. (a) and (b) In- and out-links, respectively, depicting the interactions of spatially-embedded unstable nodes with all other nodes. Link widths are proportional to link weight. (c) and (d) Distribution of in- and out-strengths, respectively, of unstable nodes compared to that of stable nodes.
@@
We now turn our focus away from cores and toward the residual nodes with unstable assignments found in the recursive significance clustering process. While we leave a comprehensive investigation into the nature of nodes with unstable community assignments to future studies, we conjecture a classification of their role in network connectivity.
@@
Overall, unstable nodes have substantially smaller in-strengths and out-strengths than the stable nodes (fig. 4c, d), corresponding to the aerodromes with fewer arrivals and departures. Looking at the origins and destinations of these comparatively few flights (fig. 4a, b), we can loosely place each unstable node into one of two functional categories: (i) a leaf node having low connectivity with a single stable node, or (ii) a bridge node with connections to multiple stable nodes belonging to different cores.
@@
The unstable-leaf nodes are unstable in the sense that they are weakly connected to one core, and as such are frequently assigned to their own singleton communities in ensemble partitions. It could reasonably be argued that assigning these leaf nodes to their adjacent cores (i.e., the core to which the node they are linked to belongs) is a more apt depiction of mesoscale system function. This is not to diminish the value of marking them as unstable, however, as one should be cautious of relying on these individual nodes to be jointly assigned with other nodes of their adjacent cores in community structure because of this low connectivity.
@@
On the other hand, unstable-bridge nodes have community assignments that are somewhat independent from that of their adjacent cores, leading to their interpretation as the uncertain (fuzzy) borders between communities. We recognize that in many real networks, a hierarchical community structure is a preferable portrayal of network organization compared to the simplified non-overlapping representation [38, 39, 40, 41]. Unstable-bridge nodes may be better described as nodes having joint membership to multiple (overlapping) communities in hierarchical structure.
V Conclusions

### v2

# Identifying robust features of community structure in complex networks

@@ Abstract
Network science has presented community detection as a valuable tool for revealing functional modules in complex systems rooted in the wiring architectures of complex networks. The varying procedures of community detection can produce, however, divisions of a network into communities that vary considerably in structure but are deemed to be of similar merit. This is especially problematic when the network is constructed on uncertain data, since small changes to the network’s configuration can cause radically different structure to be detected. To reconcile with the ambiguity in interpreting degenerate network partitions as representations of the underlying system function, we introduce a recursive significance clustering scheme that identifies the subsets of nodes having stable joint community assignments under network perturbation. These robust node groups are referred to here as cores, and represent well-supported features of network structure as distinct from the nodes with unstable community assignments. We show that cores characterize the variability inherent to non-overlapping community structure in networks and are cohesive under temporal evolution of the network.
@@ I Introduction
Characterizing complex systems is one of the great modern scientific challenges due to the difficulty of accurately and completely deriving their collective behaviors from their atomic components [1]. Enter complex networks, whose anatomies encode the intricate interactions between the discrete components of a complex system [2, 3]. Despite varying greatly in form, composition, and scope, complex networks often share common organizing principles [4].
Of particular interest in network science is the structure of a network’s wiring architecture, since structure of the network represents function of the system [1]. One manner in which many complex networks arrange themselves is into communities: groups of nodes with high internal link density that are sparsely linked with others [5]. In non-overlapping community structure, each node is assigned to exactly one community. Communities may be understood as representations of functional modules embedded within the networked system [6]. Scaling the lens of analysis to the mesoscopic interactions between communities omits superfluous detail without obscuring significant relationships in the system, providing a simplified view of interaction patterns [7].
As community detection is an NP-hard problem [8], many algorithms utilize stochastic searches [9, 10, 11, 12] to optimize an objective function that idiosyncratically defines the quality of a network partition [13, 14]. This exposes results to degenerate solutions of dissimilar community structures that are of similar merit [15, 16, 17, 18, 19]. Small changes to the network’s components (due to, e.g., measurement noise, model uncertainty, or temporal evolution) can exacerbate this issue, causing radically different structures to be detected [18, 20, 21].
@@
If a network exhibiting degenerate community structure makes the interpretation of a single partition misleading [31, 15, 32], and consensus solutions can blur relevant information captured by different partitions [20], how can results be interpreted in a scientific context? Ideally, the consistent features of the high-scoring community structures should be considered to gain insight into the structure of the network [33, 34]. We present here a recursive significance clustering scheme to reveal cores: subsets of nodes that are jointly assigned to communities with significance (e.g., fig. 1c). Cores are the features of network community structure that are well-supported by the data.
We first introduce the recursive significance clustering technique (section II) and accompanying metrics on network cores (section III). We then demonstrate the analytical capabilities of this approach by example on a real network (section IV) and validate it by studying the stability of community structure in synthetic networks (section V). All materials and methods used are described in full detail in section VII.
@@ II Recursive significance clustering
For the results of community detection to be of value, the extent to which detected communities are statistically significant as opposed to coincidental must be considered [35, 26]. After all, it is possible that a network has very weak or no community structure, meaning that detected communities are fragile and merely an artifact of the detection method used [15, 20]. A distinct optimal solution should be robust to small changes to the network’s configuration owing to the gap in quality to competing optima. In turn, the matter of community structure significance is reformulated as a question of how robust that structure is to network perturbations [26].
To identify the consistent structure in an ensemble of partitions, significance clustering procedures seek to distinguish between stable and unstable node assignments to communities [36, 20]. Stable or significant nodes of a community are flagged based on the proportion of ensemble partitions in which they are jointly assigned to a community. This approach discovers only the largest significant subset of nodes belonging to a community, but there may exist other mutually exclusive subsets of nodes in that community that are jointly assigned across the solution ensemble, but not with the largest or other subsets. This limitation betrays another: a (single) reference network partition is used to identify these significant subsets, necessarily dictating the quantity and delineation of significant subsets found. As a single partition neglects relevant information available in the whole solution landscape, this reliance on a reference partition as the mould for significant subsets is undesirable.
We define the robust cores of a network as the subsets of nodes that are jointly assigned to communities, capturing the consistent features of the landscape of community structures. For significance level $\alpha$, we require that all nodes of a core are jointly assigned in a $(1-\alpha)$-fraction of partitions to relax the requirement. The arrangement of cores into communities may vary among solutions, but that of nodes into cores is robust; in this light, cores are nearly indivisible constituents of community structure [33]. Then, delineating cores is akin to assessing the community structure variability, as they are the invariant units that compose dissimilar partitions. Nodes that do not regularly cluster with any other nodes, and nodes belonging to cores too small relative to the size of the network to contribute meaningful insight into system function, are regarded as unstable or noisy assignments.
@@
    The empirical network has its components perturbed repeatedly to generate many replicate networks that together capture modeled variability of and/or uncertainty in the network’s configuration.
@@
    An optimization technique is used to find the largest subset of nodes that are all assigned to the same community in a $(1-\alpha)$-fraction of partitions; these nodes are designated as a core, removed from the node pool, and then the next largest core is searched for, and so on until all cores of meaningful size are extracted.
This recursive approach has the advantage of operating without a reference partition; instead, all detected community structures are intrinsically used to discover the well-supported features of network structure. Section VII provides a detailed account of how we implement each step.
@@ III Viewing community structure variability
We have established cores as the elemental units of nodes that (i) are arranged like building blocks to form communities and (ii) are nearly indivisible. Identifying exactly how these cores coalesce and how stable they are as a unit provides a direct view of structural patterns in a degenerate community structure landscape and, thus, illuminates the robust features of network structure.
For this, we define the coalescence frequency of a union of cores to be the proportion of partitions in which the nodes of those cores are all assigned to the same community and without other cores. Notably, the coalescence frequency of one core is equal to the proportion of partitions where that core does not merge with other cores, instead consistently forming a community by itself.
In addition, the stability of a core is the proportion of partitions in which the nodes of the core are jointly assigned to a community. By definition, it is guaranteed to be within the interval $[1-\alpha,1]$ for significance level $\alpha$. The sum of every coalescence frequency of core unions including some core will equal that core’s stability.
@@ IV Application
To illustrate this approach, we analyze the cores of a real network. We begin by introducing the network and the recursive significance clustering setup (section IV.1). We then examine how the network cores reveal robust structure in the network (section IV.2) and how this structure is cohesive over its temporal evolution (section IV.3), and briefly examine the nature of nodes with unstable community assignments (section IV.4).
@@ IV.1 Alaska–Alaska flight network
Flight networks benefit from community detection to uncover the structure of flight routes [37, 38]. The network of flights in Alaska (AK), United States, serves as an interesting case study due to the key role that flight plays in the region’s transportation system [39].
We consider three temporal snapshots of the AK–AK flight network: calendar years 2017, 2020, and 2023. In the 2017 (resp. 2020; 2023) network, there are 3,578 (3,285; 3,616) directed links weighted by the number of aggregate flights in that year from one of the 334 (328; 330) airports to another, for 401,153 (275,890; 364,285) total annual flights. (See section VII.1 for information on data compilation). These years were selected to study flight traffic before, during, and after the largest impacts of the COVID-19 pandemic, a major disruption in aviation activity [40], which affords a prime opportunity to analyze the temporal behaviour of robust network cores.
We assume flights are independent events and resample each link’s weight from a Poisson distribution with rate parameter equal to that link’s weight in the empirical network to form 1000 replicate networks [36]. We then find the robust cores of the network’s community structure, detected using Infomap [11], at a significance level of 0.05, accepting only cores of at least five nodes. This is completed for each temporal snapshot. (See section VII for a detailed description of this procedure.)
@@ IV.2 Reckoning with dissimilar structure
Figure 2: Cores of the 2017 temporal snapshot of the AK–AK flight network. (a) Spatially embedded cores. Nodes with unstable community assignments are shaded grey. (b) Core coalescence frequency and stability.
Spatially embedded cores of the 2017 AK–AK flight network snapshot with their coalescence frequencies and stability are shown in fig. 2 (find that of the 2020 and 2023 snapshots in appendix A). For visualization, we modify the UpSet plot [41] to show the coalescence frequency for any combination of cores in tandem with the stability of each core. To manage possible combinatorial explosion, only core combinations with a coalescence frequency of at least 0.05 are displayed.
From fig. 2, we can appreciate that the AK–AK flight network has very consistent community structure, with each core forming its own community much more frequently than not. When cores do unite as communities, it always involves the large (by geographic span and node count) core 1. Speaking to the system, core 1 includes the Ted Stevens Anchorage International Airport (IATA: ANC; marked with a star), which brokers a tremendous volume of AK flight traffic, making it a highly influential node in the network as measured by virtually any centrality index. This is responsible for the vast geographic distribution of core 1 compared to other cores: the connectivity of ANC recruits airports to core 1 that otherwise have no strong local connection structure. The coalescence frequencies involving core 1 show that it sometimes merges with cores 2, 3, 4, 6, and 7. We imagine this as it irregularly absorbing the smaller cores in close proximity to it, indicating some presence of rather coarse community structure engulfing the west coast and much of the inland in the ensemble.
@@ IV.3 Temporal cohesion of cores
Now we examine the cores of successive network snapshots to assess structural changes in the network over time. An alluvial diagram [36] shows this in fig. 3. Streamlines show the core assignments of nodes changing between snapshots, coloured by their initial core assignment to track them in time. Since the node set of the network is not static in time, the flow of a node assignment is not shown if the node is not present in a contiguous temporal snapshot.
Figure 3: Core assignment changes of nodes through the 2017, 2020, and 2023 temporal snapshots of the AK–AK flight network. Nodes with unstable community assignments are placed in the grey-shaded block.
It is immediately apparent that each core is largely cohesive in time, i.e., cores continue to be stable through successive temporal snapshots. Following the streamlines, we see that some nodes change their stability with flight volume. Otherwise, the only core dynamics are the absorbing of cores 3 and 8 to core 1 in 2020 — with core 3 splitting shortly thereafter in 2023 — and core 1 ejecting a new core in 2020, which then rejoins core 1 in 2023.
An anomalous change in flight volume between an individual airport and ANC can obfuscate or reveal small-scale substructure due to the influence of ANC in the network. An increase in flight volume from 2017 occurred between ANC and Beluga Airport (BVU) of core 9 to catalyze the merge with core 1 in 2020, and a decrease between ANC and Homer Airport (HOM) for the emergence of the smaller core split from core 1. The core size threshold could be adjusted post hoc to exclude these small cores if their inclusion is felt to be more distracting than informative.
@@ IV.4 Speculating on the nature of unstable nodes
Figure 4: Wiring architectures and strength distributions of nodes with unstable community assignments for the 2017 AK–AK flight network. (a) and (b) In- and out-links, respectively, depicting the interactions of spatially embedded unstable nodes with all other nodes. Link widths are proportional to link weight. (c) and (d) Distribution of in- and out-strengths, respectively, of unstable nodes compared to that of stable nodes.
@@
We now turn our focus away from cores and toward the residual nodes with unstable assignments found in the recursive significance clustering process. While we leave a comprehensive investigation into the nature of nodes with unstable community assignments to future studies, we conjecture a classification of their role in network connectivity based on observations in the study of the AK–AK flight network.
@@
On average, unstable nodes have substantially smaller in- and out-strengths than the stable nodes (fig. 4c, d), corresponding to aerodromes with fewer arrivals and departures. Looking at the origins and destinations of these comparatively few flights (fig. 4a, b), we can loosely place each unstable node into one of two functional categories [42]: (i) a leaf node having low connectivity with a single stable node, or (ii) a branch node with connections to multiple stable nodes belonging to different cores.
@@
The unstable-leaf nodes are unstable in the sense that they are weakly connected to one core and, as such, are frequently assigned to their own singleton communities in ensemble partitions. It could reasonably be argued that assigning these leaf nodes to their adjacent cores (i.e., the core to which the node they are linked to belongs) is a more apt depiction of mesoscale system function. However, this is not to diminish the value of marking them as unstable, as one should be cautious of relying on these individual nodes to be jointly assigned with other nodes of their adjacent cores because of this low connectivity.
@@
On the other hand, unstable-branch nodes have community assignments that are somewhat independent from that of their adjacent cores, leading to their interpretation as the uncertain (fuzzy) borders between communities. We recognize that in many real networks, a hierarchical community structure is a preferable portrayal of network organization compared to the simplified non-overlapping representation [43, 44, 45, 46]. Unstable-branch nodes may be better described as nodes having joint membership to multiple (overlapping) communities in hierarchical structure.
V Community structure stability
While the AK–AK flight network demonstrated very strong community structure, it is important to characterize the cores found — and not found — in networks with weak or no community structure. The strength of the community structure in real networks can range from strong (few intercommunity links) to weak (many intercommunity links) [47]. We make use of the Lancichinetti–Fortunato–Radicchi (LFR) benchmark network [48], which is generated in the likeness of real networks, exhibiting many of the same common properties, including planted community structure. We pay particular attention to the mixing parameter $\mu$. Since each node shares a $(1-\mu)$-fraction of its links with its own community and a $\mu$-fraction with the rest of the network, $\mu\in[0,1]$ can be considered a slider with which to tune the strength of the community structure.
Ideally, we should find a large proportion of nodes in the network to be unstable when $\mu$ is large, and most nodes stable at small $\mu$. This would reflect the lack of clear community structure when there is greater mixing between communities. Beyond the critical point $\mu=0.5$, communities are no longer defined in the strong sense, with each node sharing fewer links with its community than the rest of the network [49]. In general, the success in recovering planted communities through modularity maximization has been shown to decline steeply approaching this threshold and beyond [48, 50]. We detect communities in undirected and unweighted LFR benchmark networks by maximizing modularity with the Louvain method [9]. Since we are concentrated on the positioning of links, we perturb the network by randomly redirecting a small percentage of its links. Recursive significance clustering is carried out as in the AK–AK flight network. (See section VII.)
Figure 5: Proportion of nodes with stable community assignments in LFR benchmark networks with mixing parameter $\mu$ and average degree $\langle k\rangle$.
To summarize the content of fig. 5, LFR benchmark networks with higher link density $\langle k\rangle$ are more stable at the same level of mixing $\mu$. This implies that the community structure in a network with sparse link density is more sensitive to link rearrangement, even when the number of links redirected is proportional to this density.
At low mixing, at most one core is associated with each planted community. When community structure is defined only in the weak sense beyond $\mu=0.5$, the structure is highly unstable. Between the low and high mixing regimes is a phase transition in which increasingly more planted communities are unstable. That is, entire communities of nodes become the so-called unstable-branch nodes in the network. This trait also partially accounts for a baseline instability at low mixing, where the randomly generated wiring of one or more (comparatively small) communities is not robust. This is in addition to the presence of unstable-leaf nodes that are easily severed from their community at all levels of mixing, but more prominent in networks with sparser link density.
We point out the stability plateau that emerges around $\mu=0.25$ for the two networks with higher link density $\langle k\rangle=20,25$. This increase in stability with $\mu$ from baseline instability is an exception to the monotonic non-increasing behaviour of the curves. It could suggest that, for denser networks, some degree of mixing actually fortifies the network structure against rewiring. A study of this phenomenon is left to future work.
We see then from this experiment that the existence of cores in a network is symptomatic of robust community structure. This is in a local sense: significant community structure in one part of a network is identified independently of the significance of structure elsewhere in the network and without a reference partition prescribing the locality.
VI Conclusions